1/1

Chānglíng Wáng ,  Chānglíng Wáng ,
The green Skirt of the Girl’s  Cǎi lián qǔ

translated by Shirley Yiping Zhang

The green skirts of the girls’ seem as being cut by the lotus leaves, The pretty lotuses are blossoming to the girls’ pink faces...

Héyè luóqún yīsè cái

fúróng xiàngliǎn liǎnbiā kāi Luàrù chízhōng kànbújiàn, wéngē shǐjué yǒurēnlái.

Both of them mix into the lotus pool and the girls could not been seen at all... We just realize that some poeple are coming after hearing their beautiful songs...

Translation of the ĉina poem “Cǎi lián qǔ” by Chānglíng Wáng into English by Shirley Yiping Zhang.

Arg-908-1812 (2013-07-11 22:17:14)

Author of this ĉina poem is Chānglíng Wáng .

Arg-908-1811 (2013-07-11 23:39:45)

Tiu ĉi  poemo  troviĝas  en

http: // www. ebridge. cn/ new/ music/ shirley1. php? sno= 531 . Pri la poeto vidu la retejon http: // en. wikipedia. org/ wiki/ Wang_Changling .

http://www.poezio.net/version?poem-id=908&version-id=1812,1811